# Diffusion Denoising as a Certified Defense against Clean-label Poisoning

### v1

@@ Abstract
We present a certified defense to clean-label poisoning attacks. These attacks work by injecting a small number of poisoning samples (e.g., 1%) that contain $p$-norm bounded adversarial perturbations into the training data to induce a targeted misclassification of a test-time input. Inspired by the adversarial robustness achieved by denoised smoothing, we show how an off-the-shelf diffusion model can sanitize the tampered training data. We extensively test our defense against seven clean-label poisoning attacks and reduce their attack success to 0–16% with only a negligible drop in the test time accuracy. We compare our defense with existing countermeasures against clean-label poisoning, showing that the defense reduces the attack success the most and offers the best model utility. Our results highlight the need for future work on developing stronger clean-label attacks and using our certified yet practical defense as a strong baseline to evaluate these attacks.
@@ 1 Introduction
A common practice in machine learning is to train models on a large corpus of data. This paradigm empowers many over-parameterized deep-learning models but also makes it challenging to retain the quality of data collected from various sources. This makes deep-learning-enabled systems vulnerable to data poisoning (Rubinstein et al., 2009; Nelson et al., 2008b; Biggio et al., 2012; Jagielski et al., 2018; Shafahi et al., 2018; Carlini, 2021)—where an adversary can alter a victim model’s behaviors by injecting malicious samples (i.e., poisons) into the training data.
@@
But in practice, it may be challenging to make modifications to a training dataset used for fully-supervised classification, because they are often relatively small. As a result, recent work has developed clean-label poisoning attacks (Shafahi et al., 2018; Zhu et al., 2019; Aghakhani et al., 2021; Turner et al., 2019; Saha et al., 2020; Huang et al., 2020; Geiping et al., 2021b), where the attacker aims to control the victim model’s behavior on a specific test input by injecting a few poisons that visually appear to be correctly-labeled, but in fact include human-invisible adversarial perturbations.
@@
We introduce a defense against clean-label poisoning attacks inspired by defenses to test-time adversarial examples. Existing poisoning defenses fall into two categories certified and heuristic. Certified defenses offer provable guarantees, but often significantly decrease the utility of defended models at test-time, making them impractical (Ma et al., 2019; Levine & Feizi, 2020; Wang et al., 2022; Zhang et al., 2022). Heuristic approaches (Suciu et al., 2018; Peri et al., 2020; Hong et al., 2020; Geiping et al., 2021a; Liu et al., 2022) demonstrate their effectiveness against existing attacks in realistic scenarios. However, these defenses rely on unrealistic assumptions, such as the defender knowing the target test-time input (Suciu et al., 2018), or are evaluated against specific poisoning adversaries, leaving them improvable against adaptive attacks or future adversaries.
@@
Our contributions. We first make two seemingly distant goals closer by presenting a certified defense against clean-label poisoning that also minimizes the decrease in clean accuracy. For any $p$-norm bounded adversarial perturbations to the training data, we ensure a certified accuracy higher than the prior certified defenses. The model trained on the tampered data classifies a subset of test-time input $x$ (or $x+\delta$, where $||\delta||_{\ell_{p}}$ in clean-label backdoor poisoning) correctly. To achieve this goal, we leverage the recent diffusion probabilistic diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Nichol & Dhariwal, 2021). We use an off-the-shelf diffusion model to denoise the entire training data before training a model. In §3, we theoretically show how one achieves a certified accuracy under $2$-norm bounded adversarial perturbation to the training data.
@@
Removing adversarial perturbations in the training data before model training, we can decouple the certification process from the training algorithm. Second, we leverage this computational property our defense has and present a series of training techniques to alleviate the side-effects of employing certified defenses, i.e., the utility loss of a defended model. To our knowledge, we are the first certified defense that decouples these two processes. We minimize the utility loss by employing the warm-starting (Ash & Adams, 2020). We train a model a few epochs on the tampered data or initialize its model parameters using a pre-trained model. None of them alters the training process; thus, our defense’s provable guarantee holds. In §4, we show they attribute to a better performance compared to existing certified defenses.
@@
Third, we extensively evaluate our defense against seven clean-label poisoning attacks studied in two different scenarios: transfer-learning and training a model from-scratch. In transfer-learning, our defense completely renders the attack ineffective with a negligible accuracy drop of 0.5% in the certified radius of 0.1 in $2$-norm perturbations. We also reduce the attack success to 2–16% when a defender trains a model from scratch. We further compare our defense with six poisoning defenses in the prior work. We demonstrate more (or the same in a few cases) reduction in the attack success and less accuracy drop than those existing defenses.
We discuss research questions important in studying poisoning attacks and defenses but are far-neglected in the prior work. We suggest future work directions to answer them.
@@ 2 Preliminaries on Clean-label Poisoning
Threat model. A clean-label poisoning attacker causes a misclassification of a specific target test-time sample (${x}_{t}$, $y_{t}$) by compromising the training data $D_{tr}$ with poisoning samples $D_{p}$. If a victim trains a model $f$ on the poisoned training data $D_{ptr}\!=\!D_{tr}\cup D_{p}$, the resulting model $f^{*}_{\theta}$ is likely to misclassify the target instance to the adversarial class $y_{adv}$ while preserving the classification behavior of $f^{*}_{\theta}$ on the clean test-set $S$. The attacker crafts those poisons $({x}_{p},y_{p})$ by first taking a few base samples in the same domain (${x}_{b},y_{b}$) and then adding human-imperceptible perturbations $\delta$, carefully crafted by the attacker and also bound to $||\delta||_{\ell_{p}}\leq\epsilon$, to them while keeping their labels clean ($y_{b}=y_{p}$). A typical choice of the bound is $\ell_{\infty}$ or $\ell_{2}$.
@@
While minimizing the crafting objective $\mathcal{L}_{adv}(x_{t},y_{t};f^{*}_{\theta})$, the attacker also trains a model $f^{*}_{\theta}$ on the compromised training data, which is itself another optimization problem, formulated as follows:
@@
Existing clean-label attacks. Initial work (Shafahi et al., 2018; Zhu et al., 2019; Aghakhani et al., 2021) minimizes $\mathcal{L}_{adv}$ by crafting poisons that are close to the target in the latent representation space $g(\cdot)$. A typical choice of $g(\cdot)$ is the activation outputs from the penultimate layer of a pre-trained model $f(\cdot)$, publicly available to the attacker from the Internet. The attacks have shown effective in transfer-learning scenarios, where $f(\cdot)$ will be fine-tuned on the poisoned training data. The attacker chooses base samples from the target class $(x_{b},y_{adv})$ and craft poisons $(x_{p},y_{adv})$. During fine-tuning, $f(\cdot)$ learns to correctly classify poisons in the target’s proximity in the latent representation space and classify the target into the class $y_{adv}$ the attacker wants.
@@
Recent work focuses on making those poisoning attacks effective when $f$ is trained from scratch on the tampered training set. To do so, the attacker requires to approximate the gradients computed on $\mathcal{L}_{adv}$ that are likely to appear in any models. Huang et al. (2020) address this challenge by meta-learning; the poison-crafting process simulates all the possible initialization, intermediate models, and adversarial losses computed on those intermediate models. A follow-up work by Geiping et al. (2021b) alleviates the computational overhead of the prior work by proposing an approximation technique, gradient matching, that aligns the gradients from poisoning samples with those computed on a target.
@@
Defenses against clean-label poisoning. Early work on poisoning defenses focuses on filtering out poisons from the training data. Nelson et al. (2008a) and Suciu et al. (2018) compute the training samples negatively impacting the classification results on targets. But in practice, the defender does not know which test-time samples are the attack targets. Follow-up work (Rubinstein et al., 2009; Peri et al., 2020; Tran et al., 2018) leverages unique statistical properties (e.g., spectral signatures) that can distinguish poisons from the clean data. All these approaches depend on the data they use to compute the properties. Recent work (Geiping et al., 2021a; Borgnia et al., 2021; Hong et al., 2020; Liu et al., 2022) thus reduces the dependency by proposing data- or model-agnostic defenses, adapting robust training or differentially-private training. While shown effective against existing attacks, those defenses are not provable.
@@
A separate line of work addresses this issue by proposing certifiable defenses (Levine & Feizi, 2020; Wang et al., 2022; Weber et al., 2023), which guarantee the correct classification of test-time samples when an adversary compromises the training data with the $k$ poisons whose perturbation radius of $r$. Levine & Feizi (2020); Wang et al. (2022)’s approach is majority voting, where the defender splits the training data into multiple disjoint subsets, train models on them, and run majority voting of test inputs over those models. BagFlip (Zhang et al., 2022) proposes a model-agnostic defense that uses the idea of randomized smoothing (Cohen et al., 2019) studied in adversarial robustness. However, the certification offered by these works has only shown effective in toy datasets like MNIST, and it has been under-studied whether they will scale to practical scenarios, e.g., CIFAR10 or TinyImageNet. The computational demands of these defenses further hinder their deployment in practice.
@@ 3 Diffusion Denoising as Certified Defense
We aim to certify the prediction of a model $f_{\theta}$ trained on a poisoned training data $D_{ptr}$. To formally define this goal, we first define the perturbation bound $\epsilon$ as the $\ell_{p}$ distance between two datasets, computed by taking the sum across $\ell_{p}$ perturbations of all images. The perturbation space $P^{\pi}_{\epsilon}$ is the set of all the datasets $D_{ptr}$ obtained by tampering samples with the crafting algorithm $\pi$ with the bound $\epsilon$. Operating within this perturbation space $P$, for any given test-time sample $(x_{t},y_{t})$, can produce a certificate that
$\mathbf{Pr}_{{D_{ptr}}\in P^{\pi}_{\epsilon}(D_{tr})}\left[f_{\theta\leftarrow D_{ptr}}(x_{t})=f_{\theta\leftarrow D_{tr}}(x_{t})\right]>1-\alpha$
with arbitrarily high probability $1-\alpha$, we guarantee that training on any poisoned dataset $D_{ptr}\in P^{\pi}_{r}(D_{tr})$ in the perturbation bound $\epsilon$ will classify the example $\mathbf{x}_{t}$ the same way as it was classified on the non-poisoned dataset $D_{tr}$.
@@ 3.1 Robustness Guarantee
Intuition. Randomized smoothing (Cohen et al., 2019) guarantees that a fixed classifier $f_{\theta}$ will classify any adversarial example $x^{\prime}$ correctly with high probability. To make this argument, it first shows that even for extremely large values of noise $\delta$, the example $x+\delta$ will be classified correctly. Thus, the analysis of randomized smoothing is able to show that a (much smaller) worst-case direction $\delta_{adv}$ will not change the prediction of the classifier.
Here we make a similar argument, but change where the noise is added. Given a dataset $D_{ptr}$, where some examples may be slightly perturbed to poison the model, we want to guarantee that any model $f_{\theta}$ trained on $D_{ptr}$ will classify the test examples correctly. To achieve this, we train many models $f^{i}_{\theta}$ on datasets from $P^{\pi}_{\delta}$ (with extremely large perturbations $\delta$); if each of these classifiers consistently labels the test data correctly, then we can be guaranteed that (much smaller) worst-case poisoning directions $\delta_{adv}$ can not change the prediction of the classifier.
Randomized smoothing shows that adversarial noise to the test example $x$ will not cause misprediction because even for large quantities of Gaussian noise $x+\delta$, the predictions remain consistent. In our setting, adversarial noise to the model’s training dataset $D_{tr}$ will not cause misprediction because even for large quantities of Gaussian noise $\delta$ to the training set the predictions of $f_{\theta\leftarrow D_{ptr}}$ remain consistent.
Details. Standard randomized smoothing transforms a non-robust classifier $f$ into a smoothed (robust) classifier $g$:
$g(x)=\arg\max_{c\in\mathcal{Y}}\mathbf{Pr}_{\delta\sim N(0,\sigma^{2}I)}[f_{\theta}(x+\delta)=c].$
But clean-label poisoning does not add any test-time perturbations to the target samples. Instead, it alters the training dataset, which then causes the models trained on the dataset to be different. Therefore, for our use-case we define:
$g(x)=\arg\max_{c\in\mathcal{Y}}\mathbf{Pr}_{\theta\sim\text{train}(D_{ptr}\sim P^{\pi}_{\delta})}[f_{\theta}(x)=c].$
Here, each model $f_{\theta}$ is obtained by training a new classifier on a perturbed version of the training dataset.
3.2 (Diffusion) Denoising for the Robustness
Diffusion denoising for the robustness. We utilize off-the-shelf DDPMs (Sohl-Dickstein et al., 2015; Ho et al., 2020; Nichol & Dhariwal, 2021) to denoise adversarial perturbations added to the training data and as a result, provide the robustness to clean-label poisoning. A naive adaptation of randomized smoothing to our scenario is to train multiple models on the training data with Gaussian noise augmentation or using adversarial training. But we need to add noise to the data that does not make it look natural or they are computationally demanding. We thus avoid using these approaches to train models, instead we want to remove the (potentially compromised) training set before training. Note that the robustness guarantee that randomized smoothing offers holds regardless of how a model is trained. We can still certify the robustness against clean-label poisoning. The fact that we use the denoising may make the guarantee a bit loose, but we show in our evaluation that the loose guarantee is sufficient to render existing attacks ineffective. Our denoising process is shown in Algorithm 1.
Algorithm 1 Noise, denoise, train, classify
1:  function DenoiseTrainClassify($f,\sigma,D,x,n$)
@@
12:   return $t^{*},\alpha_{t^{*}}$
Algorithm 2 Randomized smoothing (Cohen et al., 2019)
1:  function Predict($f,\sigma,$ $D$$,x,n,\alpha$)
2:   counts $\leftarrow$ DenoiseTrainClassify($f,\sigma,D,x,n_{0}$)
@@
5:   if BinomPValue($n_{A},n_{A}+n_{B},0.5\leq\alpha$) return $\hat{c_{A}}$
6:   else return ABSTAIN
@@
8:  function Certify($f,\sigma,$ $D$$,x,n_{0},n,\alpha$)
9:   counts0 $\leftarrow$ DenoiseTrainClassify($f,\sigma,D,x,n_{0}$)
@@
11:   counts $\leftarrow$ DenoiseTrainClassify($f,\sigma,D,x,n$)
12:   $p_{A}\leftarrow$ LowerConfBound(counts[$\hat{c_{A}}$], n, $1-\alpha$)
13:   if $p_{A}>1/2$ return $\hat{c_{A}}$ and radius $\sigma\Phi^{-1}(p_{A})$
14:   else return ABSTAIN

### v2

# Certified Robustness to Clean-Label Poisoning Using Diffusion Denoising

@@ Abstract
We present a certified defense to clean-label poisoning attacks under the $\ell_{2}$-norm. These attacks work by injecting a small number of poisoning samples (e.g., 1%) that contain bounded adversarial perturbations into the training data to induce a targeted misclassification of a test input. Inspired by the adversarial robustness achieved by randomized smoothing, we show how an off-the-shelf diffusion denoising model can sanitize the tampered training data. We extensively test our defense against seven clean-label poisoning attacks in both $\ell_{2}$ and $\ell_{\infty}$-norms and reduce their attack success to 0–16% with only a negligible drop in the test accuracy. We compare our defense with existing countermeasures against clean-label poisoning, showing that the defense reduces the attack success the most and offers the best model utility. Our results highlight the need for future work on developing stronger clean-label attacks and using our certified yet practical defense as a strong baseline to evaluate these attacks.
@@ 1 Introduction
A common practice in machine learning is to train models on a large corpus of data. This paradigm empowers many over-parameterized deep-learning models but also makes it challenging to retain the quality of data collected from various sources. This makes deep-learning-enabled systems vulnerable to data poisoning (Rubinstein et al., 2009; Nelson et al., 2008b; Biggio et al., 2012; Jagielski et al., 2018; Shafahi et al., 2018; Carlini, 2021)—where an adversary can alter a victim model’s behaviors by injecting malicious samples (i.e., poisoning samples) into the training data.
@@
But in practice, it may be challenging to make modifications to entire training data used for supervised classification and may be easy to detect label corruptions. As a result, recent work has developed clean-label poisoning attacks (Shafahi et al., 2018; Zhu et al., 2019; Aghakhani et al., 2021; Turner et al., 2019; Saha et al., 2020; Huang et al., 2020; Geiping et al., 2021b), where the attacker aims to control the victim model’s behavior on a specific test input by injecting a few poisons that visually appear to be correctly-labeled, but in fact include human-invisible adversarial perturbations.
@@
We present a certified defense against clean-label poisoning attacks in $\ell_{2}$-norm inspired by defenses to adversarial examples. Existing poisoning defenses fall into two categories certified and heuristic. Certified defenses offer provable guarantees, but often significantly decrease the utility of defended models at inference, making them impractical (Ma et al., 2019; Levine & Feizi, 2020; Wang et al., 2022; Zhang et al., 2022). Heuristic approaches (Suciu et al., 2018; Peri et al., 2020; Hong et al., 2020; Geiping et al., 2021a; Liu et al., 2022) demonstrate their effectiveness against existing attacks in realistic scenarios. However, these defenses rely on unrealistic assumptions, such as the defender knowing the target test input (Suciu et al., 2018), or are evaluated against specific poisoning adversaries, leaving them ineffective against adaptive attacks or future adversaries.
@@
Our contributions. First, we make two seemingly distant goals closer: presenting a certified defense against clean-label poisoning that also minimizes the decrease in clean accuracy. For $\ell_{2}$-norm bounded adversarial perturbations to the training data, we either achieve a certified accuracy higher than existing ones or require less computations to attain a comparable certified accuracy. The model trained on the tampered data classifies a subset of test input $x$ (or $x+\delta$, where $||\delta||_{\ell_{2}}$ in clean-label backdoor poisoning) correctly. To achieve this goal, we leverage the recent denoising diffusion probabilistic models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Nichol & Dhariwal, 2021). We use an off-the-shelf diffusion model to smooth the entire data before training a model and provide certified predictions.
@@
Removing adversarial perturbations in the training data before model training, we can decouple the certification process from the training algorithm. Second, we leverage this computational property our defense has and present a series of training techniques to alleviate the side-effects of employing certified defenses, i.e., the utility loss of a defended model. To our knowledge, we are the first certified defense that decouples these two processes. We minimize the utility loss by employing the warm-starting (Ash & Adams, 2020). We train a model a few epochs on the tampered data or initialize its model parameters using a pre-trained model. None of them alters the training process; thus, our defense’s provable guarantee holds. In §4, we show they attribute to a better performance compared to standard training.
@@
Third, we extensively evaluate our defense against seven clean-label poisoning attacks studied in two different scenarios: transfer-learning and training a model from-scratch. In transfer-learning, our defense completely renders the attack ineffective with a negligible accuracy drop of 0.5% in the certified radius of 0.1 in $\ell_{2}$-norm perturbations. We also reduce the attack success to 2–16% when a defender trains a model from scratch. We further compare our defense with five poisoning defenses in the prior work. We demonstrate more (or the same in a few cases) reduction in the attack success and less accuracy drop than those existing defenses.
@@ 2 Preliminaries on Clean-label Poisoning
Threat model. A clean-label poisoning attacker causes a misclassification of a specific target test sample (${x}_{t}$, $y_{t}$) by compromising the training data $D$ with poisoning samples $D_{p}$. If a victim trains a model $f$ on the poisoned training data $D^{\prime}\!=\!D\cup D_{p}$, the resulting model $f^{*}_{\theta}$ is likely to misclassify the target instance to the adversarial class $y_{adv}$ while preserving the classification behavior of $f^{*}_{\theta}$ on the clean test-set $S$. The attacker crafts those poisons $({x}_{p},y_{p})$ by first taking a few base samples in the same domain (${x}_{b},y_{b}$) and then adding human-imperceptible perturbations $\delta$, carefully crafted by the attacker and also bound to $||\delta||_{\ell_{p}}\leq\epsilon$, to them while keeping their labels clean ($y_{b}=y_{p}$).
@@
While minimizing the crafting objective, the attacker also trains a model $f^{*}_{\theta}$ on the tampered training data, which is itself another optimization problem, formulated as follows:
@@
Existing clean-label attacks. Initial work (Shafahi et al., 2018; Zhu et al., 2019; Aghakhani et al., 2021) minimizes $\mathcal{L}_{adv}$ by crafting poisons that are close to the target in the latent representation space $g(\cdot)$. A typical choice of $g(\cdot)$ is the activation outputs from the penultimate layer of a pre-trained model $f(\cdot)$. The attacks have shown effective in transfer-learning scenarios, where $f(\cdot)$ will be fine-tuned on the tampered training data. The attacker chooses base samples from the target class $(x_{b},y_{adv})$ and craft poisons $(x_{p},y_{adv})$. During fine-tuning, $f(\cdot)$ learns to correctly classify poisons in the target’s proximity in the latent representation space and classify the target into the class $y_{adv}$ the attacker wants.
@@
Recent work make those poisoning attacks effective when $f$ is trained from scratch on $D^{\prime}$. To do so, the attacker requires to approximate the gradients computed on $\mathcal{L}_{adv}$ that are likely to appear in any models. Huang et al. (2020) employs meta-learning; the poison-crafting process simulates all the possible initialization, intermediate models, and adversarial losses computed on those intermediate models. Geiping et al. (2021b) reduces this computational overhead by proposing gradient matching, that aligns the gradients from poisoning samples with those computed on a target.
@@
Defenses against clean-label poisoning. Early work on poisoning defenses (Nelson et al., 2008a; Suciu et al., 2018) focuses on filtering out poisons from the training data. Follow-up work (Rubinstein et al., 2009; Peri et al., 2020; Tran et al., 2018) leverages unique statistical properties (e.g., spectral signatures) that can distinguish poisons from the clean data. All these approaches depend on the data they use to compute the properties. Recent work (Geiping et al., 2021a; Borgnia et al., 2021; Hong et al., 2020; Liu et al., 2022) thus reduces the dependency by proposing data- or model-agnostic defenses, adapting robust training or differentially-private training. While shown effective against existing attacks, those defenses are not provable.
@@
A separate line of work proposes certified defenses (Levine & Feizi, 2020; Wang et al., 2022; Weber et al., 2023), which guarantee the correct classification of test samples when an adversary compromises the training data. Levine & Feizi (2020); Wang et al. (2022)’s approach splits the training data into multiple disjoint subsets, trains models on them, and runs majority voting of test inputs over these models. BagFlip (Zhang et al., 2022) proposes a model-agnostic defense based on randomized smoothing (Cohen et al., 2019), studied in adversarial robustness. But, the computational demands hinder their deployment in practice.
@@ 3.1 Robustness Guarantee
We formally show that our defense provides a provable robustness guarantee against clean-label poisoning attacks, particularly under $\ell_{2}$-norm perturbations. Due to space constraints, we only include important theorems and corollaries here and provide the full details in Appendix C.
Cohen et al. (2019) shows that a smoothed classifier $g$, constructed via randomized smoothing, is robust around a test input $x$ within a certified radius of $R$. The following theorem from the original study establishes this connection:
Theorem 3.1 (Robustness guarantee by Cohen et al. (2019)).
Let $f:\mathbb{R}^{d}\rightarrow\mathcal{Y}$ be any deterministic (or random) function, $\varepsilon\sim\mathcal{N}(0,\sigma^{2},I)$, and $g$ to be the smoothed classifier robust under $\varepsilon$. Suppose $y_{A}\in\mathcal{Y}$ and $p_{A},p_{B}\in[0,1]$ satisfy:
$\mathbb{P}(f(x+\varepsilon)=y_{A})\geq p_{A}>p_{B}\geq\underset{y\neq y_{A}}{\max}\mathbb{P}(f(x+\varepsilon)=c)$
then $g(x+\delta)=y_{A}$ for all $||\delta||_{2}<R$, where
$R=\frac{\sigma}{2}\big{(}\Phi^{-1}(p_{A})-\Phi^{-1}(p_{B})\big{)}$
and $\Phi$ is the inverse of the standard Gaussian CDF.
Theorem 3.1 does not impose specific restrictions on how the smoothed classifier $g$ is constructed. A common approach prior work employs is to apply additive Gaussian noise to the training data or to the internal activations (Lecuyer et al., 2019) at each mini-batch during training. But this approach is not compatible with our setting where we want to smooth the training data. To address this challenge, we adopt the approach by Weber et al. (2023), which obtains a smoothed classifier $g$ through a method that aligns with our setting:
Definition 3.2 ($g$ considered in Weber et al. (2023)).
Suppose a base classifier $h$, learned from the training data $D$ and to maximize the conditional probability distribution over labels $y$. Then the smoothed classifier $q$ is defined by:
$q(y|x,D)=\mathbb{P}_{X,D}(h(x+X,D+\mathcal{D})=y),$
where $X\sim\mathbb{P}_{X}$ is random variables, and $\mathcal{D}\sim\mathbb{P}_{\mathcal{D}}$ denotes smoothing distributions. $\mathcal{D}$ is a collection of $n$ i.i.d random variables, composed of $\mathcal{D}^{i}$s, where each $\mathcal{D}^{i}$ will be added to a training samples in $D$. $\mathbb{P}_{X}$ is the noise distribution for test inputs, and $\mathbb{P}_{\mathcal{D}}$ is the noise distribution for the training data. The final, smoothed classifier then assigns the most likely class to an instance $x$ under $q$, so that:
$g(x,D)=\operatorname*{arg\,max}_{y}q(y|x,D)$
Our work considers the smoothed classifier $g$ defined in 3.2. We then establish our robustness guarantee, based on the guarantee against backdoor attacks, which is completely different from clean-label poisoning, as follows:
Algorithm 1 Noise, denoise, train, and classify
1:  fn NDTClassify($f,\sigma,D,x,n$)
@@
12:   ret $t^{*},\alpha_{t^{*}}$
Algorithm 2 Predict and certify (Cohen et al., 2019)
1:  fn Predict($f,\sigma,$ $D$$,x,m,\alpha$)
2:   counts $\leftarrow$ NDTClassify($f,\sigma,D,x,m_{0}$)
@@
5:   if BinomPVal($n_{A},n_{A}+n_{B},0.5\leq\alpha$) ret $\hat{c_{A}}$
6:   else ret ABSTAIN
@@
8:  fn Certify($f,\sigma,$ $D$$,x,m_{0},m,\alpha$)
9:   counts0 $\leftarrow$ NDTClassify($f,\sigma,D,x,m_{0}$)
@@
11:   counts $\leftarrow$ NDTClassify($f,\sigma,D,x,m$)
12:   $p_{A}\leftarrow$ LowerCfBound(counts[$\hat{c_{A}}$], $m$, $1-\alpha$)
13:   if $p_{A}>1/2$ ret $\hat{c_{A}}$ and radius $\sigma\Phi^{-1}(p_{A})$
14:   else ret ABSTAIN
Theorem 3.3 (Robustness guarantee by Weber et al. (2023)).
Suppose $g$ is the smoothed classifier defined above with smoothing distribution $Z:=(X,\mathcal{D})$ with $X$ taking values in $\mathbb{R}^{d}$ and $\mathcal{D}$ being a collection of $n$ independent $\mathbb{R}^{d}$-valued random variables, $\mathcal{D}=(\mathcal{D}^{1},...,\mathcal{D}^{n})$. Let $\Omega_{x}\in\mathbb{R}^{d}$ and $\Delta:=(\delta^{1},\delta^{2},...,\delta^{n})$ for backdoor patterns $\delta^{i}\in\mathbb{R}^{d}$. Let $y_{A}\in\mathcal{Y}$ and $p_{A},p_{B}\in[0,1]$ such that $y_{A}=g(x,D)$ and
$q(y_{A}|x,D)\geq p_{A}>p_{B}\geq\underset{y\neq y_{A}}{\max}q(y|x,D).$
If the optimal type II errors, for testing the null $Z\sim\mathbb{P}_{0}$ against the alternative $Z+(\Omega_{x},\Delta)\sim\mathbb{P}_{1}$, satisfy
$\beta^{*}(1-p_{Z};\mathbb{P}_{0},\mathbb{P}_{1})+\beta^{*}(p_{B};\mathbb{P}_{0},\mathbb{P}_{1})>1,$
then $y_{A}=\operatorname*{arg\,max}_{y}q(y|x+\Omega_{x},D+\Delta)$.
Theorem 3.3 establishes that the smoothed classifier $q$, as defined in 3.2, trained on data containing perturbations $\Delta$, is robust around a test input $x$ with the perturbation $\Omega_{x}$.
Corollary 3.4 (Certified radius by Weber et al. (2023)).
Let $\Delta=(\delta_{1},\delta_{2},...,\delta_{n}$), $\Omega_{x}$ be $\mathbb{R}^{d}$-valued backdoor patterns, and $D$ be a training dataset. Suppose that for each $i$, the smoothing noise on the training features is $\mathcal{D}^{i}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\sigma^{2}\mathbbm{1}_{d})$. Let $y_{A}\in\mathcal{Y}$ such that $y_{A}=q(x+\Omega_{x},D+\Delta)$ with class probabilities satisfying:
$q(y_{A}|x+\Omega_{x},D+\Delta)$ $\geq p_{A}$
$>p_{B}\geq\underset{y\neq y_{A}}{\max}q(y|x+\Omega_{x},D+\Delta)$
If the perturbations on the training samples are bounded by
$\sqrt{\sum_{i=1}^{n}||\delta_{i}||_{2}^{2}}<\frac{\sigma}{2}\big{(}\Phi^{-1}(p_{A})-\Phi^{-1}(p_{B})\big{)},$
it is guaranteed $y_{A}=q(x+\Omega_{x},D)=q(x+\Omega_{x},D+\Delta)$.
Now we can simplify Corollary 3.4 by assuming an adversary compromises $r\leq n$ training samples with the largest noise pattern’s $\ell_{2}$-norm bound, $||\delta||_{2}$:
$||\delta||_{2}<\frac{\sigma}{2\sqrt{r}}(\Phi^{-1}(p_{A})-\Phi^{-1}(p_{B})\big{)},$
A typical choice of $r$ in clean-label poisoning is $\sim$10%. But when computing our certificate bound in §3, we consider the worst-case adversary capable of compromising the entire training set. In this case, $r\!=\!n$. Under this formulation, we observe that against the strongest adversary who can compromise the entire training data $n$, the robustness bound is much smaller compared to that against the realistic attacker capable of compromising only 10% of the training data.
Our certification process trains $m$ models that return the prediction of $x$, the target test sample, as $y_{A}$. Then the lower bound of $p_{A}$ becomes $\alpha^{1/m}$, with the probability of at least $1-\alpha$. Plugging this into the above equation yields:
$||\delta||_{2}<\frac{\sigma}{\sqrt{m}}(\Phi^{-1}(p_{A})\big{)}$
Corollary 3.5 (Our robustness guarantee against clean-label poisoning.).
Consider the smoothed classifier $q$ from Theorem 3.3, with no test input perturbation $X$ and the smoothing mechanism being diffusion denoising, which applies a Gaussian process to add and remove noise from the input. Under this setting, it provides the same robustness guarantee for clean-label poisoning attacks.
Proof.
By setting $X\equiv 0$ in Theorem 3.3 (and therefore $\Omega_{x}=0$), the smoothed classifier will output the same label prediction $y_{A}$ within the same $\ell_{2}$-bound for the perturbations applied to the training data of size $n$. ∎
3.2 (Diffusion) Denoising for the Robustness
Diffusion denoising for the robustness. We utilize off-the-shelf DDPMs (Sohl-Dickstein et al., 2015; Ho et al., 2020; Nichol & Dhariwal, 2021) to denoise adversarial perturbations added to the training data and as a result, provide the robustness to clean-label poisoning. A naive adaptation of randomized smoothing to our scenario is to train multiple models on the training data with Gaussian noise augmentation or using adversarial training. But we need to add noise to the data that does not make it look natural or they are computationally demanding. We thus avoid using these approaches to train models, instead we want to remove the (potentially compromised) training set before training. We use a single-step denoising, as shown in Pseudocode 1.
Figure 1: Certified radius and accuracy attained by denoising the CIFAR10 training data with varying $\sigma$ values in {0.1, 0.25, 0.5, 1.0} (left) and by adding Gaussian random noise to the training data with $\sigma$ values in {0.25, 1.0} (right).